OpenStudy (anonymous):

I need help to check this equation for extraneous solutions. 4|3x+4|=4x+8 I need the steps to if possible

OpenStudy (debbieg):

Once you find your solutions, just plug each back in and make sure that it is a true equality. If it isn't true for some solution, then that's an extraneous solution.

OpenStudy (anonymous):

Yes i know that but i don't know what to do with that 4 on the far left outside of the the absolute value equation.

OpenStudy (debbieg):

Isolate the absolute value part first. So just divide both sides by 4.

OpenStudy (anonymous):

Okay thank you I thought that was how.

OpenStudy (debbieg):

Sure thing. :)

OpenStudy (anonymous):

\[\left| 3x+4 \right|=3x+4 if 3x+4 \ge 0 or x \ge \frac{ -4 }{ 3 }\] \[\left| 3x+4 \right|=-\left( 3x+4 \right), if 3x+4 < 0, x <\frac{ -4 }{ 3 }\]